OpenStudy (anonymous):

The spring in a toy pop-gun has a spring constant of 85 N/m. To push the cork out of the barrel, the spring is compressed by 0.030 m. Part 1: Calculate the elastic potential energy of the cork. Part 2: Find the speed of the 0.002 kg cork as is leaves the barrel of the toy pop-gun. **Assume no losses from friction. How would I solve and show the work for this? :/ Thank you! :)

OpenStudy (michele_laino):

hint: the formula which gives the requested potential energy is: \[\Large EP = \frac{1}{2}k{\delta ^2}\] where k= 85 N/m, and \delta=0.03 meters

OpenStudy (michele_laino):

in order to answer ot part B, we have to keep in mind that all the potential energy will transform into kinetic energy of the cork, namely: \[\Large EP = KE = \frac{1}{2}m{v^2}\] where m is the mass of the cork, and v is the requested velocity. EP is the potential energy computed in part A.. So you have to solve this equation: \[\Large v = \sqrt {\frac{{2EP}}{m}} = ...?\]

OpenStudy (michele_laino):

oops.. all the potential energy will transform itself into kinetic energy

OpenStudy (anonymous):

so for part 1, i get 1/2(85)(0.03) = 1/2(2.55) = 1.275 ? for the potential energy? and for part 2, i am confused as to what to plug in? @Michele_Laino

OpenStudy (michele_laino):

part 1 the right computation is: \[\Large PE = \frac{1}{2} \times 85 \times {\left( {0.03} \right)^2} = \frac{{85 \times 9}}{{200}} = 3.825\;Joules\]

OpenStudy (michele_laino):

so, we can write: \[\Large v = \sqrt {\frac{{2PE}}{m}} = \sqrt {\frac{{2 \times 3.825}}{{0.002}}} = ...?\]

OpenStudy (anonymous):

ohh okay! i forgot to square it in part 1! so 3.825 J would be the final solution? for part 1? that is the potential energy of the cork? and then for part 2, calculated, v=61.84658438? so the final solution for part 2 is the speed (velocity) is 61.85? not sure about the units though :/ would it be m/s ? :/ @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!! thank you!! :)

OpenStudy (michele_laino):

thank you!! :)